Last weekend Joel traveled by train to visit his grandparents. On Friday he boarded an express train that traveled at an average rate of 80 miles per hour. On Sunday he returned home on a slower train that traveled at an average rate of 48 miles per hour. The total time Joel spent on the trains last weekend was 8 hours. How many hours did Joel spend on the train on Friday?
A.5 hr
B.2 hr
C.3 hr
D.8 hr

Answer by ikleyn(53765)

.
Last weekend Joel traveled by train to visit his grandparents. On Friday he boarded an express train that traveled at an average rate
of 80 miles per hour. On Sunday he returned home on a slower train that traveled at an average rate of 48 miles per hour.
The total time Joel spent on the trains last weekend was 8 hours. How many hours did Joel spend on the train on Friday?
~~~~~~~~~~~~~~~~~~~

The both way distances are the same, so the time spent is inversely proportional to the speeds.


The ratio of the speed "to there" to the speed "back" is 80/48 = 10/6 = 5/3.  


Therefore, the ratio of times spent is 3/5 counting time "to there" to the time "back".


It makes the answer clear and obvious: the time "to there" was 3 hours, while the time "back" was 5 hours.

Answer. Joel spent 3 hours on Friday. Answer C).